OpenStudy (anonymous):

find the distance between (2,5) and the midpoint of (-4,1) and (2,5) would I use the rule ..(x1-x2)+(y1-y2)

OpenStudy (anonymous):

Distance - \[d = \sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\] Mid point - \[ \left( \frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2 }\right)\]